OpenStudy (abmon98):

if log10x=a and log10y=b express log10sqrt(10x/y^3) in terms of a and b

OpenStudy (abmon98):

\[\log_{10} x=a \]\[\log_{10} y=b\]\[x=10^a\]\[y=10^b\] \[\log_{10} \sqrt{10*10^a/(10^b)^3}\]

ganeshie8 (ganeshie8):

simplify using exponent rules

OpenStudy (abmon98):

\[\log_{10} \sqrt{100a}/\sqrt{1000b^3}\]

OpenStudy (abmon98):

Should we multiply the denominator by sqrt of 1000B^3

OpenStudy (abmon98):

\[\log_{10} \sqrt{100^a*1000^*(-3b)}\]

ganeshie8 (ganeshie8):

\[\log_{10} \sqrt{10*10^a/(10^b)^3}\]

ganeshie8 (ganeshie8):

\[\large \log_{10} \sqrt{10^{a+1}/10^{3b}}\]

OpenStudy (abmon98):

I dont Understand why is 10^a+1 and 10^3b instead of 100^a and 1000^3b

ganeshie8 (ganeshie8):

thats a good question, so you're asking why \(10*10^{a}\) becomes \(10^{a+1}\) right ?

OpenStudy (abmon98):

yes

ganeshie8 (ganeshie8):

lets try this : \(10^3 = 10 *10*10\) \(10^4 = 10 *10*10 * 10\) \(\cdots \) \(10^a = 10*10*10* \cdots a~ times\) right ?

ganeshie8 (ganeshie8):

if we multiply 10 again, clearly we need to increase the exponent : \(10^{a+1} = 10 * (10*10*10*\cdots a~times )\) \(~~~~~~~~~~ = 10* 10^a \)

ganeshie8 (ganeshie8):

see if that looks more or less convincing

ganeshie8 (ganeshie8):

We can generalize this rule : \(\large x^m * y^n = x^{m+n} \)

ganeshie8 (ganeshie8):

\(\large 10*10^a\) \(\large 10^1*10^a\) \(\large 10^{a+1}\)

OpenStudy (abmon98):

iam convinced now by the last part thank you so much @ganeshie8 :)

ganeshie8 (ganeshie8):

good to hear :)